Systems Staff Meeting Notes 6/4/01

Systems Staff Meeting Notes
June 4, 2001

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Fluke Meter Training – Geoffrey has opted out, following the review of the training attended by Bobbie and Will. It was more like an infomercial. Bill notes that the basic training for it was purchased and it should be given here.
Review Minutes of 5/21/01
  1. Vault doors – no news, still awaiting word. Did receive word from Facilities that the posts in the tape vault are necessary for structural support and no more should be removed.
  2. SACS lectures – Vision series – Bobbie and Will have signed up to attend this Thursday & Friday.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. May be doing some more ordering this month – streaming video server is a possibility.
  2. Printers ordered last week – one big color printer for Resource Services and four small HP’s for RS (2), MSL (1) and Special Collections (1).
  1. Great browser meltdown – 2 weeks ago. NT3 – primary domain controller blue screened. It was the central repository for the master browsers – LMHost files added to the confusion. NT3 was rebuilt, all of the servers had to be rebooted. Printing was lost, Xerox printing was lost – problem was finally fixed on Memorial Day. PDC was removed from the LMHosts file - also killed the redundant WINS servers, should have only 10.99. NT5 is the only WINS server now. A lot of these problems should go away with W2K.
  2. SQUID problem this morning – couldn’t resolve DNS. There was no error message, just reinitialized it.
  3. Exchange attachment problem returned with a vengeance. Bill found a W2K document that pointed to resetting the archive bit during backups. We’ve changed the way that backups are done; hopefully it has fixed the problem.
  4. Office XP – we’re not considering it at the moment. MS is changing the licensing model – yearly licensing vs. one time. It would expire in October & the cost would depend on update frequency.
  5. Will is building a server to run the W2K upgrades. He’ll probably send a hyperlink to Systems staff – it takes about an hour and a half to run the upgrade from the server. He’ll upgrade his this afternoon. So far, we’ve upgraded NT & Win’98 successfully. Some Dells that have 2 video cards have caused trouble - one of the cards needs to be removed prior to the upgrade.
  6. Hoax virus – we did not get hit too bad – only one reported incident.
Software Projects Status
  1. Personnel mods are done.
  2. PAULA to SQL – is finished, will convert over tomorrow when Cynthia returns.
  3. Card swipe – waiting for new database product to arrive that allows for data storage.
  4. Library Stats – Winston is moving Jack’s old data, coding is done, but there is already one change – an extra field needs to be added.
  5. ILL Content Site – need to move on this.
DP Coordinator
  1. New student, Oleah Hodge, has replaced Laura Williams. We should now have full coverage.

Back to the Systems Homepage

Last updated June 11, 2001
by Debra Harris